OpenStudy (anonymous):

Use synthetic division to find P(2) for P(x) = x^4 + 3x^3 - 6x^2 - 10x + 8 Choices :) A.) 2 B.) 28 C.) 4 D.) -16

jimthompson5910 (jim_thompson5910):

You don't need synthetic division really. You can just plug in x = 2 and evaluate.

OpenStudy (anonymous):

what does it have to equal to ?

jimthompson5910 (jim_thompson5910):

P(x) = x^4 + 3x^3 - 6x^2 - 10x + 8 P(2) = (2)^4 + 3(2)^3 - 6(2)^2 - 10(2) + 8 P(2) = ???

OpenStudy (anonymous):

I got 76 ? that's not a choice

jimthompson5910 (jim_thompson5910):

P(2) = (2)^4 + 3(2)^3 - 6(2)^2 - 10(2) + 8 P(2) = 16 + 3(8) - 6(4) - 10(2) + 8 P(2) = 16 + 24 - 24 - 20 + 8 P(2) = 4

jimthompson5910 (jim_thompson5910):

not sure where you went wrong, but I'm guessing you lost a sign somewhere

OpenStudy (anonymous):

so you don't do the exponents ?

jimthompson5910 (jim_thompson5910):

what do you mean by "do exponents"?

OpenStudy (anonymous):

Like if you do 2^4 is 16 and 3(2) ^3 is 216 and 6(2) ^2 is 144 and 10(2) is 20 that's how I got 76 I added 16 + 216 - 144 - 20 + 8

jimthompson5910 (jim_thompson5910):

you do 3(2)^3 = 3(8) = 24 NOT 3(2)^3 = 6^3 = 216 because exponents come before multiplication in PEMDAS

OpenStudy (anonymous):

O, ok thanks so much for explaining :) it makes more since now

jimthompson5910 (jim_thompson5910):

yw